# Adaptive Canonicalization With Application to Invariant Anisotropic Geometric Networks

### v1

@@ E.2 Construction Details for Anisotropic Point Cloud Networks
We first describe adaptive canonicalization for the DeepSet architecture [Zaheer et al., 2017], which is a general framework for learning function on sets and has UAT [Wagstaff et al., 2022]. Given a point cloud represented as $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\subset{\mathbb{R}}^{3}$, DeepSet computes a permutation-invariant representation by encoding each point with an MLP $\phi$, aggregating features with summation, and applying a global function $\xi$:
$\mathbf{x}^{\prime}=\xi\left(\sum_{i=1}^{N}\phi(\mathbf{x}_{i})\right)$
For classification into $D$ classes, we assign each class $d\in\{1,\ldots,D\}$ a learned canonical form ${\mathbf{R}}_{d}^{\square}\in\mathcal{SO}(3)$ for a prior $h_{d}$ chosen to maximize class separability in one vs. rest manner:
@@
where $\Psi_{d}$ is a class-specific classifier for class $d$. Then, the class scores are computed by $s_{d}=\psi_{d}\left(\xi\left(\sum_{i=1}^{N}\phi(\mathbf{x}_{i}{\mathbf{R}}_{d}^{\square\top})\right)\right)$ for all $d\in\{1,\ldots,D\}$ , and then passed through sigmoids. The loss is trained by summing over $D$ binary cross-entropies.
@@
Next, we apply adaptive canonicalization in PointNet [Qi et al., 2017]. For a point cloud $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\in{\mathbb{R}}^{3}$, PointNet processes each point independently with a shared MLP $\phi$, and then aggregates points features using a symmetric function (e.g., max-pooling)
@@
For each class $d$ with a prior $h_{d}$, we define a class-specific orientation ${\mathbf{R}}_{d}^{\square}$ obtained by maximizing discriminative power:
@@
where $\Psi_{d}$ is the class-specific classifier conducting binary classification in the one vs. rest manner. Each ${\mathbf{R}}_{d}^{\square}$ emphasizes the characteristics most relevant to class $d$. The scoring head is denoted by $s_{d}=\Psi_{d}(\operatorname{Pool}(\phi(\mathbf{x}_{1}{\mathbf{R}}_{d}^{\square\top}),\ldots,\phi(\mathbf{x}_{N}{\mathbf{R}}_{d}^{\square\top})))$. Then we define the per-class head score by a sigmoid and use binary cross-entropy per class and sum over classes.
@@
We then consider the application of adaptive canonicalization to DGCNN [Wang et al., 2019]. Consider the input ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ as a point cloud with points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\in{\mathbb{R}}^{3}$. DGCNN constructs dynamic $k$-nearest graphs by computing $G=(V,E)$ where $E=\{(i,j):j\in\operatorname{kNN}(\mathbf{x}_{i},k)\}$. Then, the edge convolution is performed by computing the edge features and applying a max pooling:
@@
Consider a classification task with $D$ classes. For every class $d\in\{1,\ldots,D\}$, we define a class-specific orientation ${\mathbf{R}}_{d}^{\square}$ obtained by optimizing a class-dependent prior $h_{d}$:
@@
where $\Psi_{d}$ is the class-specific classifier. The canonical form ${\mathbf{R}}_{d}^{\square}$ emphasizes the characteristics most relevant to class $d$. Then, the classifier is applied to the output under ${\mathbf{R}}_{d}^{\square}$, i.e., $s_{d}=\Psi_{d}(\operatorname{Pool}_{(i,j)\in E}(\operatorname{ReLU}(\Psi((\mathbf{x}_{j}-\mathbf{x}_{i}){\mathbf{R}}^{\square\top},\mathbf{x}_{i}{\mathbf{R}}^{\square\top}))))$ and defines the per-class head score by a sigmoid. We then use binary cross-entropy per class and sum over classes.
Efficient Optimization with Random Sampling and Local Refinement.
A direct search over rotations can be implemented with quaternion [Shoemake, 1992] or Rodrigues’ formula [Rodrigues, 1840], but exhaustive exploration of all rotations is costly. For computation efficiency, we adopt a sample and refine scheme. Specifically, to optimize each class’s canonical rotation efficiently, we first sample a small set of random rotation matrices to serve as initial candidates. For each candidate, we compute the prior-weighted objective and then perform the refinement by taking the gradient descent steps. Note that this inner loop of gradient descent is not differentiated through: the refined rotation is treated as a fixed input to the subsequent forward pass and backpropagation. In this way, we avoid the high cost of full meta-optimization over all rotations. This keeps the expensive optimization out of backpropagation while providing a good canonical orientation per class.
Adaptive canonicalization operates like a max-pooling step: the network evaluates multiple candidate orientations and selects the one that maximizes the prior. The chosen orientation is passed forward, and during backpropagation, gradients flow only through this selected orientation.
@@ F.1 Illustrative Toy Problems: Grid Signal Orientation Tasks
We use a three-layer network with a hidden feature dimension chosen from $\{32,64,128\}$ and ReLU activation functions. The learning rate is selected from $\{10^{-3},10^{-4},10^{-5}\}$. Batch size 100. All models are implemented in PyTorch and optimized with the Adam optimizer [Kingma and Ba, 2014]. Experiments are conducted on an Nvidia DGX A100. The output of the GNN is then passed to an MLP, followed by a softmax classifier.
@@ F.2 Graph Classification on TUDataset
We consider five graph classification benchmarks from TUDataset [Morris et al., 2020]: MUTAG, PTC, ENZYMES, PROTEINS, and NCI1. The dataset statistic is reported in Tab. LABEL:tab:statistics. Following the random split protocol [Ma et al., 2019, Ying et al., 2018, Zhang et al., 2019b], we partition the dataset into 80% training, 10% validation, and 10% testing. Results are averaged over 10 random splits, with mean accuracy and standard deviation reported.
@@
The hidden dimension is set to be 128. The models are implemented using PyTorch, optimized with the Adam optimizer [Kingma and Ba, 2014]. An early stopping strategy is applied, where training halts if the validation loss does not improve for 100 consecutive epochs. The hyperparameters are selected through a grid search, conducted via Optuna [Akiba et al., 2019], with with the learning rate and weight decay explored in the set $\{1e^{-2},1e^{-3},1e^{-4}\}$, the pooling ratio varying within $[0.1,0.9]$ with step $0.1$, and the number of layers ranging from 2 to 9 in a step size of 1. The output representations are then passed into an MLP followed by a softmax layer, and predictions are obtained by optimizing a cross-entropy loss function. Experiments are conducted on an Nvidia DGX A100.
@@ F.3 Molecular Classification on OGB Datasets
We evaluate on larger-scale benchmarks from the Open Graph Benchmark (OGB) dataset [Hu et al., 2020] for classification tasks, including ogbg-molhiv, ogbg-molpcba, and ogbg-ppa. Dataset statistics are summarized in Tab. LABEL:tab:statistics The evaluation settings are followed by the OGB protocol [Hu et al., 2020].
@@
The models are implemented in PyTorch and optimized with the Adam optimizer, with training capped at a maximum of 1000 epochs and controlled by an early stopping criterion. The hidden dimension is selected from the set $\{128,256,512\}$, while the number of layers varies from 2 to 10 in steps of 1. Dropout rates are explored within the range $0,0.1,\dots,0.5]$, the learning rate is tuned within the interval $[0.0001,0.001]$, and the warmup is set as 5 or 10. Additionally, the batch size is chosen from $\{32,64,128,256\}$ and the weight decay is chosen from $\{10^{-4},10^{-5},10^{-6}\}$. All hyperparameters are tuned using Optuna [Akiba et al., 2019]. The experiments are conducted on an NVIDIA A100 GPU.
@@ F.4 ModelNet40 Point Cloud Classification
Our evaluation for point cloud classification was carried out on the ModelNet40 dataset [Wu et al., 2015], which consists of 40 object categories and a total of 12,311 3D models. Following prior studies [Wang et al., 2019, Deng et al., 2021], we allocated 9,843 models for training and 2,468 models for testing in the classification task. For each model, 1,024 points were uniformly sampled from its mesh surface, using only the xyzxyz coordinates of the sampled points. We apply on-the-fly rotation augmentation during training, following Esteves et al. [2018], Deng et al. [2021], such that the dataset size remains unchanged. At test time, each example is rotated by an arbitrary rotation. Note that on-the-fly augmentation essentially changes the training data distribution during the learning process. The purpose of comparing under rotation protocols is to assess a model’s invariance to rotational changes.
@@
For point cloud classification tasks, we compare our anisotropic geometric method with point cloud approaches, including PointNet and DGCNN architectures, as well as equivariant models based on the vector neuron framework, i.e., VN-PointNet and VN-DGCNN. We further include canonicalization baselines, CN-PointNet and CN-DGCNN, and traditional augmentation baselines in which the training set is expanded with pre-generated rotations (PointNet-Aug, DGCNN-Aug) with a factor of five ($\times 5$). The experimental results of PointNet, DGCNN, VN-PointNet, and VN-DGCNN are taken from Wang et al. [2019], Deng et al. [2021].

### v2

@@ E.4 Construction Details for Anisotropic Point Cloud Networks
We begin by briefly recalling DeepSet [zaheer2017deep], which is a general framework for learning functions on sets and has UAT [wagstaff2022universal]. Given a point cloud represented as $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\subset{\mathbb{R}}^{3}$, DeepSet computes a permutation-invariant representation by encoding each point with an MLP $\phi$, aggregating features with summation, and applying a global function $\xi$:
$\mathbf{x}^{\prime}=\xi\left(\sum_{i=1}^{N}\phi(\mathbf{x}_{i})\right).$
To apply adaptive canonicalization to DeepSet, we define, for each class $d\in\{1,\ldots,D\}$, a class-dependent canonical rotation by
@@
where $\Psi_{d}$ is the one-vs-rest classifier for class $d$, and $h_{d}(x)=x$ is the prior. Once ${\mathbf{R}}_{d}^{\square}$ is obtained, the class scores are computed as $s_{d}=\Psi_{d}(\xi(\textstyle\sum_{i=1}^{N}\phi(\mathbf{x}_{i}{\mathbf{R}}_{d}^{\square\top})))$ for all $d\in\{1,\ldots,D\}$, and passed through sigmoids to obtain class-wise probabilities. The network is then trained using the sum of $D$ binary cross-entropy losses.
@@
Given a point cloud $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\in{\mathbb{R}}^{3}$, PointNet [qi2017pointnet] processes each point independently with a shared MLP $\phi$, and then aggregates points features using a symmetric function (e.g., max-pooling)
@@
To equip PointNet with adaptive canonicalization, we define a class-specific orientation ${\mathbf{R}}_{d}^{\square}$ for each class $d$ obtained by
@@
where $h_{d}(x)=x$ is the class-$d$ prior and $\Psi_{d}$ is the class-specific classifier conducting binary classification in the one vs. rest manner. Once having ${\mathbf{R}}_{d}^{\square}$, the scoring head is denoted by $s_{d}=\Psi_{d}(\operatorname{Pool}(\phi(\mathbf{x}_{1}{\mathbf{R}}_{d}^{\square\top}),\ldots,\phi(\mathbf{x}_{N}{\mathbf{R}}_{d}^{\square\top})))$. Then we define the per-class head score by a sigmoid and use binary cross-entropy per class and sum over classes as the training loss.
@@
Consider the input ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ as a point cloud with points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\in{\mathbb{R}}^{3}$. DGCNN [wang2019dynamic] constructs dynamic $k$-nearest graphs by computing $G=(V,E)$ where $E=\{(i,j):j\in\operatorname{kNN}(\mathbf{x}_{i},k)\}$. Then, the edge convolution is performed by computing the edge features and applying a max pooling:
@@
Applying adaptive canonicalization to the DGCNN architecture, we define a class-specific orientation ${\mathbf{R}}_{d}^{\square}$ obtained by optimizing a class-dependent prior $h_{d}(x)=x$:
@@
where $\Psi_{d}$ is the class-$d$ classifier. Then, the classifier is applied to the output under ${\mathbf{R}}_{d}^{\square}$ for computing $s_{d}=\Psi_{d}(\operatorname{Pool}_{(i,j)\in E}(\operatorname{ReLU}(\Psi((\mathbf{x}_{j}-\mathbf{x}_{i}){\mathbf{R}}^{\square\top},\mathbf{x}_{i}{\mathbf{R}}^{\square\top}))))$. The per-class head score is then computed by a sigmoid, and the sum over $D$ binary cross-entropies is used as the training loss.
E.5 Truncation Canonicalization
We introduce truncation-based prior maximization on images as an additional application of our adaptive canonicalization framework. Intuitively, many image classes are invariant under removing uninformative regions: as long as the object of interest remains in the field of view, the class label should not change. We consider, in this case, the truncation as our “symmetry” and apply prior maximization over this family of transformations to select a canonical truncation. We note that this transformation family is not based on a group action.
We model images as elements of a set of $\mathcal{G}$ in the following way. Given $N\in\mathbb{N}$, consider the regular grid on the unit square $0,1]^{2}$ given by the pixels $Q_{i,j}=[\frac{i}{N},\frac{i+1}{N})\times[\frac{j}{N},\frac{j+1}{N}))$ for $0\leq i,j<N$. We define $\mathcal{G}$ to be the set of all functions $g\in L_{2}({\mathbb{R}}^{2})$ such that $\operatorname{supp}(g)\subset[0,1]^{2}$ and $g$ is piecewise constant on the grid, i.e., there exist coefficients $v_{i,j}\in[0,1]$ such that $g(x)=v_{i,j}$ for all $x\in\{Q_{i,j}\}$. Since each image $g\in\mathcal{G}$ is uniquely determined by its collection of pixel values $(v_{i,j})_{0\leq i,j<N}\in[0,1]^{N^{2}}$, we can identify (via an isometric isomorphism) $\mathcal{G}$ with with the box $0,1]^{N^{2}}\subset{\mathbb{R}}^{N^{2}}$. Since closed and bounded subsets of finite-dimensional Euclidean spaces are compact, the box $0,1]^{N^{2}}\subset{\mathbb{R}}^{N^{2}}$ is compact. Because $\mathcal{G}$ is isometrically isomorphic to $0,1]^{N^{2}}$, the set $\mathcal{G}$ is also compact (and therefore locally compact).
The truncation is parameterized by four coordinates collected in a vector $u=(x_{\mathrm{L}},y_{\mathrm{T}},x_{\mathrm{R}},y_{\mathrm{B}})\in\mathcal{U}\subset[0,1]^{4}$, where we impose the constraints $0\leq x_{\mathrm{L}}\leq x_{\mathrm{R}}\leq 1$ and $0\leq y_{\mathrm{T}}\leq y_{\mathrm{B}}\leq 1$. We consider the standard Euclidean metric in $\mathcal{U}$. Since $\mathcal{U}$ is a closed and bounded subset of $0,1]^{4}$, it is compact.
We now define the truncation transformation. For a parameter $u=(x_{\mathrm{L}},y_{\mathrm{T}},x_{\mathrm{R}},y_{\mathrm{B}})\in\mathcal{U}$, we view $u$ as encoding the top-left and bottom-right corners of an axis-aligned rectangle: $R(u)=[x_{\mathrm{L}},x_{\mathrm{R}}]\times[y_{\mathrm{T}},y_{\mathrm{B}}]\subset[0,1]^{2}$. As $u$ varies in the compact box $\mathcal{U}$, this truncation window moves continuously inside the domain containing all images. Given an image $g\in\mathcal{G}$, we define the truncation by
[…]
After truncation, the function lies in $L_{2}({\mathbb{R}}^{2})$ but not in $\mathcal{G}$ in general. We then project it back onto the finite-dimensional subspace $\mathcal{G}$ of piecewise constant functions by the orthogonal projection $P:L_{2}({\mathbb{R}}^{2})\rightarrow\mathcal{G}$ defined by
$(Pg)(x)=\sum_{i,j=1}^{N}\bar{g}_{ij}\,\mathbbm{1}_{Q_{ij}}(x),\qquad\bar{g}_{ij}=N^{2}\int_{Q_{ij}}g(y)dy.$
We then define $\kappa_{u}(g):=P(g\mathbbm{1}_{R(u)})\in\mathcal{G}$.
We now verify that this truncation family satisfies our assumptions. By construction, $\mathcal{U}$ is compact, and the image space $\mathcal{G}$ is compact. In our setting we have $\mathcal{K}=\mathcal{G}$, so $\mathcal{K}$ is also compact. The map
$\kappa_{(\cdot)}(\cdot\cdot):\mathcal{U}\times\mathcal{G}\rightarrow\mathcal{K},\quad(u,g)\mapsto\kappa_{u}(g)\in\mathcal{K}$
is continuous, and hence, by compactness of the product $\mathcal{U}\times\mathcal{G}$, it is uniformly continuous. This implies that $\{g\mapsto f(\kappa_{u}(g))\}_{u\in\mathcal{U}}$ is equicontinuous. Thus, the truncation transformation family satisfies the conditions for continuous canonicalization.
We focus on the image classification task with $D$ classes and take axis-aligned truncations as the transformation family. Let ${\mathbf{X}}\in{\mathbb{R}}^{N\times N}$ be an image. A truncation is specified by a side-length $s\in[s_{\text{min}},s_{\text{max}}]\subset[0.5,1.0]$ and a discrete top-left corner. For a given $s$, we set the truncation side length to $M(s)=\lfloor sN\rfloor$, and require the window to lie inside the image domain, i.e., $1\leq i_{0}\leq N-M(s)+1$ and $1\leq j_{0}\leq N-M(s)+1$. The corresponding axis-aligned window is $\{(i,j)\in\{1,\ldots,N\}^{2}\,\big|\,i_{0}\leq i\leq i_{0}+M(s)-1,\;j_{0}\leq j\leq j_{0}+M(s)-1\}$. The truncation then keeps only the pixels inside the window and zeros out the rest, which yields a truncated patch $C_{(i_{0},j_{0},s)}({\mathbf{X}})\in{\mathbb{R}}^{M(s)\times M(s)}$. To keep the input dimension fixed, we then rescale this cropped patch back to the original resolution $N\times N$ using a standard interpolation scheme. We denote this resizing operator by $R_{s}:\mathbb{R}^{M(s)\times M(s)}\to\mathbb{R}^{N\times N}$. The truncation operator is therefore defined as the composition $T_{(i_{0},j_{0},s)}({\mathbf{X}})=R_{s}(C_{(i_{0},j_{0},s)}({\mathbf{X}}))\in{\mathbb{R}}^{N\times N}$. The transformation family used in our method is
$\mathcal{T}=\{T_{(i_{0},j_{0},s)}\;|\;s\in[s_{\min},s_{\max}],\;(i_{0},j_{0})\text{ admissible as above}\},$
and in practice random truncations are obtained by sampling $s$ and $(i_{0},j_{0})$ from a suitable distribution under these constraints.
We introduce the truncation canonicalization as follows. Consider a symmetryless neural network $\Psi:{\mathbb{R}}^{N\times N}\rightarrow{\mathbb{R}}^{D}$ that operates on the images. We denote $\Psi=(\Psi_{d})_{d=1}^{D}$ where $\Psi_{d}:{\mathbb{R}}^{N\times N}\rightarrow{\mathbb{R}}$. The prior maximization for each class $d\in\{1,\dots,D\}$ is performed by
$T_{(i_{0},j_{0},s)}^{\square}=\underset{T_{(i_{0},j_{0},s)}\in\mathcal{T}}{\arg\max}\;h_{d}(\Psi_{d}(T_{(i_{0},j_{0},s)}({\mathbf{X}}))),$
where $h(x)=x$. Once we obtain $T_{(i_{0},j_{0},s)}^{\square}$, the class $d$ score is then computed by $\Psi_{d}(T_{(i_{0},j_{0},s)}^{\square}({\mathbf{X}}))$ are then passed through a sigmoid nonlinearity to obtain class probabilities. The training loss is the sum of $D$ binary cross-entropies.
@@ F.1 Illustrative Toy Problems: Grid Signal Orientation Tasks
We use a three-layer network with a hidden feature dimension chosen from $\{32,64,128\}$ and ReLU activation functions. The learning rate is selected from $\{10^{-3},10^{-4},10^{-5}\}$. Batch size 100. All models are implemented in PyTorch and optimized with the Adam optimizer [kingma2014adam]. Experiments are conducted on an Nvidia DGX A100. The output of the GNN is then passed to an MLP, followed by a softmax classifier.
@@ F.2 Graph Classification on TUDataset
We consider five graph classification benchmarks from TUDataset [morris2020tudataset]: MUTAG, PTC, ENZYMES, PROTEINS, and NCI1. The dataset statistic is reported in Tab. LABEL:tab:statistics. Following the random split protocol [ma2019graph, ying2018hierarchical, zhang2019hierarchical], we partition the dataset into 80% training, 10% validation, and 10% testing. Results are averaged over 10 random splits, with mean accuracy and standard deviation reported.
@@
The hidden dimension is set to be 128. The models are implemented using PyTorch, optimized with the Adam optimizer [kingma2014adam]. An early stopping strategy is applied, where training halts if the validation loss does not improve for 100 consecutive epochs. The hyperparameters are selected through a grid search, conducted via Optuna [akiba2019optuna], with with the learning rate and weight decay explored in the set $\{1e^{-2},1e^{-3},1e^{-4}\}$, the pooling ratio varying within $[0.1,0.9]$ with step $0.1$, and the number of layers ranging from 2 to 9 in a step size of 1. The output representations are then passed into an MLP followed by a softmax layer, and predictions are obtained by optimizing a cross-entropy loss function. Experiments are conducted on an Nvidia DGX A100.
@@ F.3 Molecular Classification on OGB Datasets
We evaluate on larger-scale benchmarks from the Open Graph Benchmark (OGB) dataset [hu2020open] for classification tasks, including ogbg-molhiv, ogbg-molpcba, and ogbg-ppa. Dataset statistics are summarized in Tab. LABEL:tab:statistics The evaluation settings are followed by the OGB protocol [hu2020open].
@@
The models are implemented in PyTorch and optimized with the Adam optimizer, with training capped at a maximum of 1000 epochs and controlled by an early stopping criterion. The hidden dimension is selected from the set $\{128,256,512\}$, while the number of layers varies from 2 to 10 in steps of 1. Dropout rates are explored within the range $0,0.1,\dots,0.5]$, the learning rate is tuned within the interval $[0.0001,0.001]$, and the warmup is set as 5 or 10. Additionally, the batch size is chosen from $\{32,64,128,256\}$ and the weight decay is chosen from $\{10^{-4},10^{-5},10^{-6}\}$. All hyperparameters are tuned using Optuna [akiba2019optuna]. The experiments are conducted on an NVIDIA A100 GPU.
@@ F.4 ModelNet40 Point Cloud Classification
Our evaluation for point cloud classification was carried out on the ModelNet40 dataset [wu20153d], which consists of 40 object categories and a total of 12,311 3D models. Following prior studies [wang2019dynamic, deng2021vector], we allocated 9,843 models for training and 2,468 models for testing in the classification task. For each model, 1,024 points were uniformly sampled from its mesh surface, using only the xyzxyz coordinates of the sampled points. We apply on-the-fly rotation augmentation during training, following esteves2018learning, deng2021vector, such that the dataset size remains unchanged. At test time, each example is rotated by an arbitrary rotation. Note that on-the-fly augmentation essentially changes the training data distribution during the learning process. The purpose of comparing under rotation protocols is to assess a model’s invariance to rotational changes.
@@
For point cloud classification tasks, we compare our anisotropic geometric method with point cloud approaches, including PointNet and DGCNN architectures, as well as equivariant models based on the vector neuron framework, i.e., VN-PointNet and VN-DGCNN. We further include canonicalization baselines, CN-PointNet and CN-DGCNN, and traditional augmentation baselines in which the training set is expanded with pre-generated rotations (PointNet-Aug, DGCNN-Aug) with a factor of five ($\times 5$). The experimental results of PointNet, DGCNN, VN-PointNet, and VN-DGCNN are taken from wang2019dynamic, deng2021vector.